OpenStudy (error1603):

A tuning fork produces a sound with a frequency of 1 kHz. What is the wavelength of the sound wave? Choose the nearest value from the list below. (Take the speed of sound to be 340 meters per second.) One of the answers are 680 mm 68 mm 34 mm 340 m 340 mm 34 m

OpenStudy (maths99):

Answer is 340mm

OpenStudy (error1603):

How is it 340mm? Explain.

OpenStudy (maths99):

Do you know the equation for speed, wavelength and frequency?

OpenStudy (error1603):

This is what got. λ = v/f = (340 m/s)/(1000 s^-1) = 0.340 m 0.340 m * 1000 mm/m = 340 mm

OpenStudy (maths99):

Thats the answer and the correct solution.

OpenStudy (error1603):

Oh great.

OpenStudy (maths99):

No problem